# NoMIRACL: Knowing When You Don’t Know for Robust Multilingual Retrieval-Augmented Generation

### 18 Dec 2023 v1

@@ 4.2 Baseline Model: GPT-4
Model Settings. For our experiments, we evaluate GPT-4, using the OpenAI service (API version 2023-05-15) deployed on Microsoft Azure.444learn.microsoft.com/en-us/azure/ai-services/openai/ [^5] We compared Azure API with OpenAI API across four languages in NoMIRACL and observed no noticeable difference between different GPT-4 API version providers. For all our experiments, we maintain a context window size of a maximum of 4096 tokens, a low-temperature score = 0.3 for a deterministic LLM output, and top-p sampling ratio = 0.95. We output a maximum of $50$ tokens from GPT-4.
@@
Figure 4: Vanilla zero-shot prompt template used in our work with GPT-4 OpenAI (2023) for all 18 languages in NoMIRACL. The instruction is provided in English, similar to Ahuja et al. (2023).
Figure 5: GPT-4 hallucination rate ($\mathrm{FP}/\mathrm{FP}+\mathrm{FN}$) on the non-relevant subset ($\mathrm{F}$) in NoMIRACL test split. The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ passages are judged by a human annotator as non-relevant. On average, GPT-4 OpenAI (2023) achieves a hallucination rate of 33.2% to dismiss information available in non-relevant (noisy) retrieved passages. Lower the hallucination rate is better.
Figure 6: GPT-4 error rate ($\mathrm{TN}/\mathrm{TN}+\mathrm{TP}$) on the relevant subset ($\mathrm{T}$) in NoMIRACL test split. The relevant subset contains queries with known answers, i.e., at least one of the top-$k$ ($k=10$) passages Zhang et al. (2023) judged by a human annotator is relevant. On average, GPT-4 OpenAI (2023) achieves an error rate of 14.9% to identify relevant information (relevant in answering the query) within the retrieved passages. Lower the error rate is better.
Vanilla Prompting. The choice of prompt significantly influences the performance and LLMs have been shown brittle to prompting variations, training examples or long context setups Liu et al. (2023). In this work, we evaluate GPT-4 using a zero-shot monolingual listwise prompting strategy. We construct a vanilla prompt template using all top-$k$ (oracle) passages (available in NoMIRACL) as a list of contexts along with the input query, both in the same language. We evaluate GPT-4 zero-shot, as we cannot fit few-shot exemplars due to insufficient context length. Our template provides a short description in English describing the task (Ahuja et al., 2023). Our prompt can be found in Figure 4.
Reducing Costs. Running GPT-4 is expensive. For long contexts and low-resource languages, the costs even can multiply. To limit our experimental cost, we did not exceed our prompt above 4096 tokens. To effectively fit all $k=10$ passages within the vanilla zero-shot prompt, we truncate each passage to use the first 390 tokens. Next, for evaluation, we use a sampling ratio of 50-50% (non-relevant to relevant subset) and limit GPT-4 evaluation to a maximum of 250 randomly sampled queries for all languages in both NoMIRACL relevant and non-relevant split. In our experiments, we ended up producing $\approx$10K API calls producing an expense of \$1,174 (in USD) including miscellaneous costs.
@@ 5 Experimental Results
We discuss our GPT-4 baseline results using the hallucination rate on the non-relevant subset in (§5.1) and the error rate on the relevant subset in (§5.2).
@@ 5.1 NoMIRACL Non-relevant Subset
Figure 5 shows GPT-4 hallucination rates on the NoMIRACL non-relevant subset for a maximum of 250 queries evaluated (each language) on 18 languages. On average, GPT-4 achieves a 33.2% hallucination rate, which highlights the challenge to identify all non-relevant passages in retrieved information in the prompt. The lowest hallucination rates (marked in green) are found in Swahili (sw) with 8.8% and Spanish (es) with 11.6%. Whereas, the highest (marked in red) are found in Thai (th) with 53.2% and French (fr) with 46.8%.
Further, we calculate the correlation of hallucination rates in GPT-4 with resource availability, i.e., Wikipedia corpus size in NoMIRACL. We achieve a positive correlation (Spearman’s rank correlation $\rho=0.39$, which can increase up to $\rho=0.51$ without Spanish). We hypothesize that queries in lower and medium-resource languages (smaller Wikipedia corpus) contain retrieved information that is likely to be non-relevant (easier negative), thereby making it easy for the LLM to judge as “I don’t know”. Future work can focus on improving GPT-4’s ability not to generate an answer for non-relevant knowledge. We also note that our current baseline uses a closed-box or black-box LLM; we encourage future work to explore open-source alternatives on NoMIRACL.
@@ 5.2 NoMIRACL Relevant Subset
Figure 6 shows the GPT-4 error rates on NoMIRACL relevant subset for a maximum of 250 queries (each language) on 18 languages. On average, GPT-4 achieves a low error rate of 14.9%, which is 20.3% lower than the hallucination rate, indicating a stronger ability to identify relevant passages in retrieved information. GPT-4 observes the lowest error rates with below 5% for high-resource languages such as Chinese (zh), Thai (th), Japanese (ja) or English (en). The highest error rate is observed in Telugu (te) with 42.0%. Overall, the low error rates indicate that model is able to identify relevant passages present within non-relevant passages, with a relatively higher error rate visible for low-resource languages, for example, Telugu (te), Hindi (hi) and Swahili (sw). Spanish (es) is a unique exception in the high-resource language category, where GPT-4 dismiss information confidently as “I don’t know” with a 22.0% error rate on the relevant subset.
6 Ongoing Work
Identifying LLM reasoning capabilities and robustness against first-stage retrieval errors in RAG applications has only begun. The first step in our robustness evaluation is to release the NoMIRACL dataset with both the development and test splits for all languages and provide a GPT-4 baseline to highlight the challenge in LLM robustness and facilitate more robust and accurate LLM-generated outputs in the future.
We are actively working on benchmarking a wider variety of open-source multilingual LLMs, a few notable ones are LLAMA-2 Touvron et al. (2023), FLAN-T5 Chung et al. (2022) and PolyLM Wei et al. (2023) among others. We also wish to evaluate whether better prompting strategies or fine-tuning LLMs can help improve robustness by evaluating methods such as Chain-of-Thought (CoT; Wei et al., 2022), Chain-of-Verification (CoVe; Dhuliawala et al., 2023) or recently introduced Chain-of-Noting (CoN; Yu et al., 2023).
@@ 8 Conclusion
Retrieval-augmented generation setups are effective in the factual grounding of LLM-generated output with information available in external knowledge bases. However, in cases where retrieved information does not assist in providing the answer, a robust LLM should potentially identify them. In our work, we provide NoMIRACL for evaluating LLM robustness as a binary classification task on 18 languages. We provide two subsets in NoMIRACL, the non-relevant subset, where queries contain all non-relevant passages, and the relevant subset, where queries contain at least a single relevant passage. We evaluate robustness using two metrics: hallucination rate and error rate.
In our work, we build a GPT-4 baseline on NoMIRACL. GPT-4 achieves a high 33.2% hallucination rate on the non-relevant subset and 14.9% error rate on the relevant NoMIRACL split, highlighting that GPT-4 finds it challenging to dismiss non-relevant passages over relevant passages in first-stage retrieved information. We open-source NoMIRACL and hope it encourages future work to improve LLM robustness.

### 04 Mar 2024 v2

@@ 4.3 Experimental Settings
We execute the generation of GPT-4 and GPT-3.5-turbo, using the OpenAI service (API version 2023-05-15) deployed on Microsoft Azure. [^5] We compared Azure API with OpenAI API across four languages in NoMIRACL and observed no noticeable difference between different GPT-4 API version providers. For a fair comparison across all LLMs, we maintain a maximum input sequence length of 4096 tokens. We set a low-temperature score = 0.1 for a deterministic LLM output, and a top-p sampling ratio = 0.95. We output a maximum of $50$ tokens.
@@
Figure 4: Vanilla zero-shot prompt template used in our experiments for LLM hallucination evaluation for all 18 languages in NoMIRACL. The instruction is provided in English, similar to Ahuja et al. (2023).
Figure 5: Hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN}$) on the non-relevant subset in NoMIRACL test split. The non-relevant subset contains queries with no known answers, i.e., all top-$k$ (where $k=10$) passages are judged by a human annotator as non-relevant. A majority of LLMs (except Mistral) hallucinate on the non-relevant subset. Lower the hallucination rate is better. The best model in each category is plotted (see Figure 8 for all models).
Figure 6: Error rate (in %) = $\mathrm{FN}/(\mathrm{FN}+\mathrm{TP}$) on the relevant subset ($\mathrm{T}$) in NoMIRACL test split. The relevant subset contains queries with known answers, i.e., at least one of the top-$k$ (where $k=10$) passages are judged by a human annotator as relevant. On average, a majority of LLMs (except Mistral and Aya-101) have a lower error rate by accurately identifying the answer. Lower the error rate is better. The best model in each category is plotted (see Figure 9 for all models).
Vanilla Prompting. The choice of prompt significantly influences the performance and LLMs have been shown brittle to prompting variations, training examples or long context setups Liu et al. (2023). In this work, we evaluate all baselines using a zero-shot monolingual listwise prompting strategy. We construct a vanilla prompt template using all top-$k$ (oracle) passages (available in NoMIRACL) as a list of contexts along with the input query, both in the same language. We evaluate all LLMs zero-shot, as we cannot fit few-shot exemplars due to insufficient context length (maximum of 4096 sequence length). Our template provides a short description in English describing the task (Ahuja et al., 2023). Our vanilla prompt template used in our experiments is shown in Figure 4.
Reducing Costs. Running GPT-4 is expensive and LLAMA-70B is rather slow at inference. For long contexts and low-resource languages, the costs even can multiply. To limit our experimental cost, we did not exceed our prompt above 4096 tokens. To effectively fit all $k=10$ passages within the vanilla zero-shot prompt, we truncate each passage to use the first 375 tokens. Next, due to budget constraints, we keep our evaluation to a maximum of 250 randomly sampled queries for all languages in both NoMIRACL relevant and non-relevant split. In our experiments, we ended up producing $\approx$20K API calls producing an expense of \$1,474 (in USD) including miscellaneous costs.
@@ 5 Experimental Results
We discuss our LLM robustness evaluation results using the hallucination rate on the non-relevant subset in (§5.1) and using the error rate on the relevant subset in (§5.2), and compare overall both the relevant and non-relevant capacities in (§5.3).
@@ 5.1 NoMIRACL Non-relevant Subset
Figure 5 shows hallucination rates on the NoMIRACL non-relevant subset for a maximum of 250 queries evaluated (each language) on all 18 languages for best LLM in each category (for all model results, please refer to Figure 8). Our findings indicate that all LLMs (except Mistral) hallucinate that an answer is present across all languages, thereby indicating their poor ability to abstain from answering. On average, the lowest hallucination rate of 11.7% is observed by Mistral-7B, followed by Mixtral-7x8B with 17.4%. GPT-4 achieves a 33.2% hallucination rate, which highlights the challenge of identifying non-relevant passages. FLAN-T5, LLAMA-2, Orca-2 and Aya-101 perform much worse on average across all languages, by achieving a hallucination rate of more than 80%. We hypothesize that LLMs perform poorly to identify non-relevant passages as they are highly similar to the query, but do not contain the exact answer.
Overall, the lowest hallucination rates are observed in Swahili and Yoruba. We hypothesize that queries in low-resource languages (smaller Wikipedia corpus) contain retrieved information that is likely to be non-relevant (easier negative), thereby making it easier for the LLM to judge as “I don’t know”. GPT-3.5 (cf. Figure 8) is observed with the highest deviation across languages with a hallucination rate as low as 25.2% on Swahili (sw) to 95.2% on Bengali (bn). Overall, a majority of LLMs are found to perform poorly on NoMIRACL non-relevant subset, indicating our dataset is very challenging in robustness evaluation for LLMs.
@@ 5.2 NoMIRACL Relevant Subset
Figure 6 shows error rates on NoMIRACL relevant subset for a maximum of 250 queries (each language) on 18 languages for best LLM in each category (for all model results, please refer to Figure 9). Our findings indicate that all LLMs (except Mistral) are able to identify the answer present within the relevant passage accurately. On average, Mistral-7B achieves the highest error rate of 74.9%. The lowest error rates are observed by FLAN-T5, LLAMA-2-70B, and GPT-4 which are lower than 16%. Overall, Aya-101, Mixtral-7x8B and LLAMA-2-7B perform worse on average by achieving more than a 40% error rate. Overall, a all LMs (except Mistral and Aya-101) perform well and do not suffer from errors in identifying answers in the relevant subset.
5.3 NoMIRACL Overall Comparison
A robust LLM should be able to identify the answer captured within retrieved passages in the relevant subset and abstain from answering when none of the retrieved passages contain the answer in the non-relevant subset. To measure performance across both dimensions in NoMIRACL, we plot the average model accuracy across both the non-relevant (x-axis) and relevant subset (y-axis) for all tested models in Figure 7. Overall, LLMs positioned in the top-right corner provide an optimal performance on both subsets. A majority of LLMs (such as LLAMA-2, Orca-2, FLAN-T5) in the top-left corner perform well on the relevant subset, however, hallucinate and struggle to perform well on the non-relevant subset, indicating their inability to accurately determine non-relevant passages.
On the other hand, Mistral LLMs do not suffer from hallucination on the non-relevant subset but instead overall observe a higher error rate (over 40%) on the relevant subset, indicating they are not able to confidently identify passages containing the relevant answer. Aya-101 is unable to perform well in either of the subsets. GPT-4 overall provides the performance tradeoff balancing a low hallucination and error rate on NoMIRACL relevant and non-relevant subsets, however can be costly to be computed at scale for inference.
Figure 7: Plot measuring average model accuracy across all languages on relevant (y-axis) and non-relevant subset (x-axis) in NoMIRACL. Performance towards the top-right corner (denoted by the arrow) is better.
6 Empirical Analysis
In this section, we conduct an empirical analysis of LLM outputs, with respect to both the samples in the non-relevant subset (containing hallucinations) and relevant subset (containing errors) where LLMs fail to recognize answers within the relevant passages. We conduct our ablation on the English (en) subset in NoMIRACL. We categorized each LLM output pattern as either positive or accurate (highlighted in green), unable to understand instruction (highlighted in orange) or either a hallucination or error (highlighted in red).
Non-relevant subset. As shown in section 6 (above), we observe an uniform distribution of the hallucination pattern of failed samples for a majority of LLMs by answering “Yes, answer is present” with or without additional explanation. LLMs such as LLAMA-2-7B and LLAMA-2-13B suffer from hallucinations by often repeating the question or instruction in their generation output. Mistral-7B interestingly always provides a rationale or explanation in their model response, whereas Aya-101, uses implicit memory heavily to directly provide an answer instead of grounding the answer from within the retrieved passages. Lastly, models such as Orca-2-7B tend to change the output generation style and often use synonyms such as “No, answer is not present” instead of “I don’t know”.
Relevant subset. As shown in section 6 (below), similar to the non-relevant subset, we observe a uniform distribution of accurate LLM responses. Interestingly, GPT-4, Orca-2-13B and FLAN-T5-XXL overall only provide a single output classification token, whereas models such as GPT-3.5 or Mixtral-8x7B provide an additional rationale or explanation. Similar to the non-relevant subset, LLAMA-2-7B and 13B models repeat the question and instruction in their output and Aya-101 overall uses implicit memory to answer queries.
OpenAI Mistral Orca-2 Aya LLAMA-2 FLAN-T5
GPT-4 GPT-3.5 8x7B 7B 13B 7B Aya-101 70B 13B 7B XXL
Empirical results on the non-relevant subset: Queries with all human-judged non-relevant passages
(i) Perfectly answers “I don’t know” 56.8% 54.8% 67.2% 1.6% 3.2% 9.2% 15.6% 0.0% 0.0% 0.8% 15.2%
(ii) “I don’t know” with explanation 0.0% 0.0% 3.6% 88.4% 0.0% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0%
(iii) Uses a synonym of “I don’t know” 0.4% 0.0% 1.6% 1.2% 6.0% 7.6% 0.0% 0.4% 0.0% 2.4% 6.4%
[…]
(v) Repeats question or instruction 0.0% 0.0% 0.0% 0.0% 0.0% 3.2% 0.0% 2.0% 64.8% 30.8% 0.0%
(vi) Conversation 0.0% 0.0% 0.4% 0.0% 0.0% 0.4% 0.0% 0.0% 0.0% 0.4% 0.0%
[…]
(viii) Uses implicit memory to answer 0.0% 0.0% 0.0% 0.0% 0.0% 8.8% 10.8% 0.0% 0.0% 0.0% 0.4%
Empirical results on the relevant subset: Queries with at least one human-judged relevant passage
(i) Perfectly answers “Yes, answer is present” 94.8% 45.2% 22.8% 5.2% 97.2% 71.2% 51.6% 1.2% 3.6% 7.2% 97.6%
(ii) "Yes, answer is present" with explanation 0.0% 46.0% 56.0% 31.6% 0.8% 11.2% 0.0% 98.8% 46.8% 63.2% 0.0%
[…]
(iv) Repeats question or instruction 0.0% 0.0% 0.4% 0.0% 0.0% 0.4% 0.0% 0.0% 48.8% 24.8% 0.0%
(v) Conversation 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.6% 0.0%
[…]
(vii) Uses implicit memory to answer 0.0% 0.0% 0.0% 2.4% 1.2% 14.4% 42.0% 0.0% 0.0% 0.4% 0.0%
Table 2: Empirical results on the complete NoMIRACL English (en) non-relevant (above) and relevant (below) subsets. The analysis is bracketed into three categories, where green category denotes an accurate response, orange denotes limitations in understanding the instruction and red denotes model hallucination or error respectively.
@@ 8 Conclusion
We introduce NoMIRACL, a human-annotated hallucination dataset for evaluating the performance of LLM robustness in recognizing no-answers as a binary classification task in 18 languages. Our multilingual dataset is human-annotated and constructed with 31 native speakers. We provide two subsets in NoMIRACL, the non-relevant subset, where queries contain all judged non-relevant passages, and the relevant subset, where queries contain at least a single relevant judged passage to measure both hallucination on the non-relevant and error on the relevant subset. Our experimental results indicate that existing LLMs are not robust, as we observe challenges in LLM robustness in either hallucination or error. Overall, GPT-4 achieves the best model and performance tradeoff across both subsets. We observed several hallucination patterns within LLMs. NoMIRACL can facilitate research in understanding to which extent LLMs tend to hallucinate, ultimately paving the way for building more effective and robust LLMs in the future.
9 Limitations
NoMIRACL is not perfect and like other datasets have limitations. We describe our limitations below and keep it as future work to continue to improve our dataset.
1. No finetuning. In our work, we focus on robustness evaluation on a wide variety of multilingual LLMs. We keep as future work to explore supervised fine-tuning LLMs to reduce hallucinations in LLMs, such as Chain-of-Verification (CoVe; Dhuliawala et al., 2023) or recently introduced Chain-of-Noting (CoN; Yu et al., 2023).
2. Human errors in dataset construction. Our dataset has been fully constructed using humans, thereby it may contain human errors. We conducted an additional quality checks on a subset of the NoMIRACL dataset to validate its question quality and relevance judgement as explained in Appendix C.
3. Evaluation setup. In our work, we evaluate whether a passage is either relevant or non-relevant passage for a given query, instead of evaluating actual answer spans. Coming up with reliable and accurate answers for a given query requires domain experts as annotators. Annotators can potentially highlight short extractive spans of answers within relevant passages, however, non-extractive queries can either contain multiple answers or a long-form answer, making it difficult to highlight a relevant answer span. Therefore, for NoMIRACL, we focus on the evaluation of top-$k$ passages as information contexts, which were easily judged for their relevancy by a data annotator.
4. Prompt Optimization. In our experiments, we evaluate models using a single vanilla prompt template. We leave it as future work, to further experiment with better and prompt optimzation techniques, such as including Chain-of-Thought (CoT; Wei et al., 2022).